40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15443     Entry time: Tue Jun 30 22:00:04 2020
Author: gautam 
Type: Update 
Category: Electronics 
Subject: Glitchy POX resurfaces 

This problem reared its ugly head again. I am inclined to believe the problem is electronic and not on the light, since the POY channels seem immune to this issue (see Attachment #1). I will investigate in the daytime tomorrow. Note that while the POX photodiode head has ~twice the transimpedance than POY (per measurement), the POY signal gets amplified by a ZHL-500-HLN amplifier before heading to the demod electronics (nominal gain is 19dB = x9). There is also some imbalance in the light level at the photodiodes I guess, because overall, the PDH fringe is ~twice as large for the Y arm as the X arm. Basically, the y-axes of the attached plot cannot be directly compared between POX and POY.

Mostly this is an annoyance - right now, the POX signal is only used for locking and dither aligning the X arm cavity, and so once that is done, the locking can proceed (as long as the other channels, e.g. REFL11, aren't glitching as well...)

Attachment 1: glitchyPOX.jpg  1.086 MB  Uploaded Tue Jun 30 23:01:33 2020  | Hide | Hide all
glitchyPOX.jpg
ELOG V3.1.3-